# Thera: Aliasing-Free Arbitrary-Scale Super-Resolution with Neural Heat Fields

### v3

@@ A.5 Initialization of Components
We have noticed during our experiments that the initialization of the components, $\mathbf{W}_{1}$ in Eq. 1, is important. Sitzmann et al. [41] made similar observations when periodic activation functions were first used for neural fields. The final distribution of frequencies $\left|\nu(\mathbf{W}_{1})\right|$ did not change much during training. Thus, we choose to initialize $\mathbf{W}_{1}$ such that
@@ Appendix B Continuous Upsampling Example
In Fig. 9, we provide a practical showcase of the continuous upsampling capabilities of our method.
Figure 9: Showcase of multiscale upsampling using Thera Pro with a RDN [54] backbone, shown with non-integer scaling factors.
@@ C.1 Further Metrics
×$\times 2$$\times 3$$\times 4$$\times 6$$\times 8$
LIIF [10] .934 .866 .804 .704 .636
LTE [24] .936 .868 .807 .707 .639
CiaoSR [7] .938 .871 .814 .718 .651
SRNO [45] .937 .873 .819 .738 .685
MSIT [56] .942 .882 .829 .751 .700
@@
Table 4: SSIM scores (higher is better) for several methods and scaling factors evaluated on the (hold-out) DIV2K validation set. We use the RDN backbone for all models. Some methods did not provide checkpoints at the time of writing, see Appendix F.
@@
Table 5: PSNR (Y channel) on common benchmark datasets for out-of-distribution scale factors, with an RDN [54] backbone. For some methods, code and/or checkpoints were not publicly avalibale, see Appendix F.
Table 4 shows SSIM metrics on the DIV2K validation set, which complement the PSNR values reported in Table 1. We use the SSIM implementation from torchmetrics [13]. We observe strong performance of Thera, although overall there is relatively little variance across all methods using this metric.
Furthermore, in Table 5 we show quantitative evaluations which are out of distribution both in terms of data (benchmark datasets) and in terms of scaling factors (above ×$\times 4$).
@@ C.2 Parameter Efficiency
In Fig. 10 we compare the number of additional parameters and PSNR values for various methods and individual upsampling factors on the DIV2K validation set.
$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$34.734.734.734.734.834.834.834.834.934.934.934.93535353535.135.135.135.135.235.235.235.2Thera AirThera PlusThera Pro(Ours)LIIFLTECUFCiaoSRCLITSRNOMSITModel Size (M parameters)PSNR (dB)×2 Scaling Factor
$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$3131313131.131.131.131.131.231.231.231.231.331.331.331.331.431.431.431.431.531.531.531.5Thera AirThera PlusThera Pro(Ours)LIIFLTECUFCiaoSRCLITSRNOMSITModel Size (M parameters)PSNR (dB)×3 Scaling Factor
$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$2929292929.129.129.129.129.229.229.229.229.329.329.329.329.429.429.429.429.529.529.529.5Thera AirThera PlusThera Pro(Ours)LIIFLTECUFCiaoSRCLITSRNOMSITModel Size (M parameters)PSNR (dB)×4 Scaling Factor
@@
Figure 10: Comparison of ASR methods for different scaling factors (×2, ×3, and ×4) on DIV2K. Thera consistently achieves better performance at lower parameter counts across all scaling factors.
Appendix D Hypernetwork Architecture
@@
In our implementation, Thera Air uses no feature refinement blocks on top of the backbone except a $1\times 1$ convolution mapping pixel-wise features into field parameters, as done in SIREN [41]. Thera Plus uses 6 ConvNeXt [29] blocks with $d=64$ followed by 7 ConvNeXt blocks with $d=96$ and 3 ConvNeXt blocks with $d=128$, prior to the final mapping layer. Projection blocks are added between blocks with different $d$, which consist of a layer normalization and a $1\times 1$ convolution operation. For Thera Pro, two SwinIR [25] blocks are used with depth 7 and 6, respectively, and 6 attention heads per layer.
@@
Appendix E Real-World Optical Zoom Data
To evaluate the effectiveness of our approach on real-world continuous optical zoom data, we conducted experiments using the COZ dataset [14]. We introduce Thera++, which combines two components: (1) Thera Plus (EDSR) trained on the COZ dataset, and (2) a lightweight spatial transformer network (STN) [21] that estimates just 6 parameters per image to correct domain-specific affine distortions. Despite its name, the STN is not a transformer network in the modern sense that employs attention mechanisms; rather, it is an image model block that explicitly allows the spatial manipulation of data within a convolutional neural network.
Thera++ addresses an important limitation of the COZ dataset. When looking at the data it becomes obvious that COZ samples are not perfectly aligned, resulting in $xy$-jitter between images. Additionally, dynamic objects like people, dust, leaves, and moving shadows appear inconsistently across images of the same scene, significantly increasing the noise level. These challenges make super-resolution particularly difficult for this dataset. However, Thera++ outperforms previous state-of-the-art methods, including LMI, across all scaling factors, highlighting its applicability under real-world imaging conditions, see Tab. 6.
@@ Appendix E Real-World Optical Zoom Data
MetaSR [18] 28.70 26.55 25.17 24.31 23.25
LIIF [10] 28.72 26.61 25.16 24.32 23.23
LTE [24] 28.67 26.55 25.15 24.37 23.26
SRNO [45] 28.73 26.59 25.15 24.31 23.25
LIT [9] 28.74 26.58 25.15 24.35 23.19
LMI [14] 28.86 26.66 25.22 24.39 23.29
@@
Table 6: Results (PSNR in dB) on the COZ test set.
Appendix F Reproducibility of Existing Methods
@@
CUF [44]. At the time of writing, there were no public code or checkpoints available for CUF. Therefore, we could not generate numbers for datasets and scaling factors not reported in the original paper. We have denoted those missing values with “—” in the tables. Furthermore, we could not create any qualitative samples using CUF.
CLIT [9]. For CLIT, at the time of writing, there is a public code, but no checkpoints. We have made a bona fide attempt to reproduce the models, but due to the cascaded training schedule and the large model size, the training process would require excessive amounts of compute: over a month using 8×8$\times$8 × Nvidia GeForce RTX 3090 GPUs. Unfortunately, the authors did not respond to our requests for the trained checkpoints used in their paper.
CiaoSR [7]. We found that in the official CiaoSR implementation border cropping prior to evaluating on DIV2K deviated slightly from all other methods. We thus adapted the evaluation code to match the competition and enable a meaningful comparison. All DIV2K numbers were re-computed with the corrected code, resulting in slight deviations from those reported in the original paper.
MSIT [56]. The numbers reported in the paper were achieved by training on a roughly 3$\times$ larger training set, comprising not only DIV2K but also Flickr2K. For a fair comparison with all other ASR methods used in our paper, we re-trained MSIT on DIV2K alone. We used the authors’ official code and configuration files and trained for two stages, each comprising 1050 epochs. For the second (RiM) training stage we used scaling factors between ×$\times 1$ and ×$\times 4$. Performance generally drops as a result of the smaller training set, in line with the ablation experiments reported for MSIT, where the models were also trained only with DIV2K.

### v4

@@ A.5 Initialization of Components
We have noticed during our experiments that the initialization of the components, $\mathbf{W}_{1}$ in Equation 1, is important. Sitzmann et al. (2020b) made similar observations when periodic activation functions were first used for neural fields. The final distribution of frequencies $\left|\nu(\mathbf{W}_{1})\right|$ did not change much during training. Thus, we choose to initialize $\mathbf{W}_{1}$ such that
@@ Appendix B Continuous Upsampling Example
In Figure 9, we provide a practical showcase of the continuous upsampling capabilities of our method.
Figure 9: Showcase of multiscale upsampling using Thera Pro with a RDN (Zhang et al., 2018) backbone, shown with non-integer scaling factors.
@@ C.1 Further Metrics
Table 4 shows SSIM metrics on the DIV2K validation set, which complement the PSNR values reported in Table 1. We use the SSIM implementation from torchmetrics (Detlefsen et al., 2022). We observe strong performance of Thera, although overall there is relatively little variance across all methods using this metric.
Table 4: SSIM scores (higher is better) for several methods and scaling factors evaluated on the (hold-out) DIV2K validation set. We use the RDN backbone for all models. Some methods did not provide checkpoints at the time of writing, see Appendix H.
@@
$\times 2$ $\times 3$ $\times 4$ $\times 6$ $\times 8$
LIIF .934 .866 .804 .704 .636
LTE .936 .868 .807 .707 .639
CiaoSR .938 .871 .814 .718 .651
SRNO .937 .873 .819 .738 .685
MSIT .942 .882 .829 .751 .700
@@
Thera Pro (std.) $\pm$ 2.0e-5 $\pm$ 4.0e-5 $\pm$ 8.9e-5 $\pm$ 9.7e-5 $\pm$ 5.5e-5
Table 5: Evaluation of GMSD (lower is better) on DIV2K for various methods.
Backbone Model In-distribution Out-of-distribution
$\times$2 $\times$3 $\times$4 $\times$6 $\times$12 $\times$18 $\times$24 $\times$30
EDSR-b. LIIF 0.0065 0.0354 0.0667 0.1192 0.1976 0.2319 0.2516 0.2645
LTE 0.0064 0.0351 0.0664 0.1187 0.1970 0.2316 0.2515 0.2645
Ciao 0.0063 0.0340 0.0641 0.1159 0.1942 0.2295 0.2497 0.2635
SRNO 0.0063 0.0344 0.0648 0.1165 0.1947 0.2296 0.2499 0.2631
MSIT 0.0061 0.0334 0.0636 0.1156 0.1944 0.2298 0.2502 0.2634
Thera Pro 0.0057 0.0314 0.0596 0.1094 0.1878 0.2242 0.2451 0.2591
Thera Pro (std.) $\pm$ 1.0e-5 $\pm$ 4.1e-5 $\pm$ 7.6e-5 $\pm$ 1.6e-5 $\pm$ 7.1e-5 $\pm$ 5.9e-5 $\pm$ 1.2e-4 $\pm$ 1.6e-4
RDN LIIF 0.0060 0.0330 0.0628 0.1145 0.1934 0.2285 0.2489 0.2622
LTE 0.0058 0.0327 0.0622 0.1135 0.1925 0.2280 0.2487 0.2621
Ciao 0.0058 0.0321 0.0609 0.1114 0.1897 0.2258 0.2470 0.2614
SRNO 0.0057 0.0321 0.0611 0.1118 0.1906 0.2266 0.2477 0.2613
MSIT 0.0057 0.0321 0.0610 0.1122 0.1920 0.2288 0.2501 0.2641
Thera Pro 0.0055 0.0309 0.0588 0.1083 0.1869 0.2235 0.2446 0.2588
Thera Pro (std.) $\pm$ 7.9e-6 $\pm$ 1.4e-5 $\pm$ 4.0e-5 $\pm$ 3.1e-5 $\pm$ 3.7e-5 $\pm$ 2.0e-5 $\pm$ 6.3e-5 $\pm$ 9.0e-5
Table 6: PSNR (Y channel) on common benchmark datasets for out-of-distribution scale factors, with an RDN (Zhang et al., 2018) backbone. For some methods, code and/or checkpoints were not publicly available, see Appendix H.
@@
Table 5 reports Gradient Magnitude Similarity Deviation (Xue et al., 2013) (GMSD) metrics for various methods. Thera reaches significantly lower GMSD for all backbone-scale combinations, indicating more faithful gradient structure. Note that some methods did not provide code/checkpoints and can therefore not be re-evaluated (see Section H).
Furthermore, in Table 6 we show quantitative evaluations which are out of distribution both in terms of data (benchmark datasets) and in terms of scaling factors (above $\times 4$).
@@ C.2 Parameter Efficiency
In Figure 10 we compare the number of additional parameters and PSNR values for various methods and individual upsampling factors on the DIV2K validation set.
[…]
@@
Figure 10: Comparison of ASR methods for different scaling factors ($\times$2, $\times$3, and $\times$4) on DIV2K. Thera consistently achieves better performance at lower parameter counts across all scaling factors.
C.3 Error Bars
Table 7 reports standard deviations for PSNR observed over N=3N\!=\!3 training runs that were initialized with different random seeds, complementing the main table. For SSIM and GMSD, we have reported standard deviations alongside the respective metrics in Tables 4 and 5. Standard deviations are in all cases significantly lower than performance improvements over competing methods.
Table 7: Standard deviation of PSNR (in dB) on the DIV2K validation set over N=3N\!=\!3 runs.
Backbone Method In-distribution Out-of-distribution
$\times$2 $\times$3 $\times$4 $\times$6 $\times$12 $\times$18 $\times$24 $\times$30
EDSR-b. Thera Air 0.0137 0.0143 0.0149 0.0107 0.0084 0.0055 0.0040 0.0040
Thera Plus 0.0131 0.0122 0.0114 0.0094 0.0075 0.0049 0.0060 0.0022
Thera Pro 0.0063 0.0088 0.0062 0.0034 0.0054 0.0049 0.0056 0.0034
RDN Thera Air 0.0187 0.0152 0.0113 0.0045 0.0036 0.0054 0.0022 0.0034
Thera Plus 0.0045 0.0018 0.0011 0.0011 0.0009 0.0013 0.0026 0.0024
Thera Pro 0.0025 0.0009 0.0038 0.0024 0.0021 0.0036 0.0021 0.0012
Appendix D Analysis of Learned Components & Kappa
Figure 11 shows a statistical analysis of the frequency components learned by Thera-Pro with an RDN backbone for the DIV2K training data. Components are distributed uniformly across directions, with fewer components at low frequencies and progressively more components at higher frequencies. This provides more representational capacity for the reconstruction of fine details, analogous to the typical distribution of frequency components for other decompositions such as the Fourier or discrete cosine transforms.
To investigate whether these components form a generalizable basis, we fit heat fields to images from various datasets other than the training data: Urban100, Manga109, and a highly out-of-distribution subset from the fastMRI (Zbontar et al., 2018) medical imaging dataset (single-coil knee validation split, with intensities scaled to [0,1]). In all experiments, we fix the frequency bank to the one shown in Figure 11 and optimize only the scale and shift parameters with the AdamW optimizer for 600 iterations per image, with learning rate 0.001. We use local fields that cover patches of $N\times N$ pixels, with $N\in\{3,4,6,12,18\}$. Table 8 shows that the pre-trained components reconstruct all datasets with negligible error at smaller local field sizes (in-distribution scales), and still achieve very low errors at larger field sizes up to $18\times 18$ pixels (out-of-distribution scales).
These numbers are not surprising, as the frequency bank was optimized to fit local fields and can act as an over-parametrized dictionary for smaller field sizes (e.g., $4\times 4$ GT pixels per field). Notably, it still works relatively well for large OOD sampling factors. Also, there is no obvious difference between the values for natural images and those for MRI images, which suggests that the frequency bank is a general-purpose basis, similar to Fourier or DCT components.
Table 8: Reconstruction MSE when fitting heat field grids of various sizes to out-of-distribution datasets.
Local field size $\times3$ $\times4$ $\times6$ $\times12$ $\times18$
Set5 2e-10 6e-11 1e-11 1.37e-6 2.16e-4
Urban100 4e-9 9e-10 2e-10 2.83e-6 7.54e-4
Manga109 2e-8 4e-9 7e-10 1.41e-6 3.01e-4
fastMRI 3e-11 1e-11 1e-11 8.83e-6 9.23e-4
Furthermore, Table 9 reports final, converged values of the thermal diffusivity coefficient $\kappa$. We observe that $\kappa$ is very similar across runs ($\sigma\approx 0.003$), but deviates from the theoretically derived value (for a Gaussian downsampling model, Equation 3) by a factor of $\approx 1/2$. This deviation suggests that the cutoff frequency of the anti-aliasing used in the cubic Mitchell-Netravali filter – used to downsample images during training – does not exactly match the Gaussian one used in the theoretical derivations, being more lenient towards aliasing than the theoretical model (anti-aliasing filters are tuned empirically to balance aliasing against loss of details). The result indicates that Thera is indeed able to tune $\kappa$ as necessary to approximate the downsampling characteristics seen in the training data, in a repeatable manner.
Table 9: Converged values of $\kappa$ for multiple runs.
Backbone Model Converged $\kappa$
EDSR-b. Thera Air 0.0295
Thera Plus 0.0336
Thera Pro 0.0266
RDN Thera Air 0.0300
Thera Plus 0.0360
Thera Pro 0.0293
Theoretical Gaussian 0.0702
Figure 11: Statistical distribution of the converged frequency bank of the Thera Pro run with an RDN backbone. Left: Polar scatter plot of spatial frequency and angular direction of components. Right: Corresponding marginal distributions.
Appendix E Hypernetwork Architecture
@@
In our implementation, Thera Air uses no feature refinement blocks on top of the backbone except a $1\times 1$ convolution mapping pixel-wise features into field parameters, as done in SIREN (Sitzmann et al., 2020b). Thera Plus uses 6 ConvNeXt (Liu et al., 2022) blocks with $d=64$ followed by 7 ConvNeXt blocks with $d=96$ and 3 ConvNeXt blocks with $d=128$, prior to the final mapping layer. Projection blocks are added between blocks with different $d$, which consist of a layer normalization and a $1\times 1$ convolution operation. For Thera Pro, two windowed transformer-based blocks (Liang et al., 2021; Liu et al., 2021) blocks are used with 7 and 6 layers respectively, and 6 attention heads in each layer.
@@
Appendix F Computational Complexity
Table 10 reports inference time and VRAM requirements of different Thera variants as well as comparison methods, on the $48\times 48$ pixels standard input patch size and scaling factor 4 (i.e., $192\times 192$ output size). Each Thera variant improves compute time and memory efficiency compared to methods with similar or higher parameter counts, with particularly pronounced gains over Transformer-based competitors (e.g., Thera Pro uses less than $1/4$ of the time and $1/13$ of the VRAM footprint of MSIT, at similar parameter count). All tests were performed on an NVIDIA GeForce RTX 3090 Ti GPU.
Table 10: Comparison of runtime and VRAM footprint of various methods with an EDSR-baseline backbone.
Method Add. params. Time VRAM
MetaSR 0.45 M 3.72 ms 834 MiB
LIIF 0.35 M 13.8 ms 604 MiB
LTE 0.49 M 16.7 ms 574 MiB
CiaoSR 1.43 M 37.5 ms 3894 MiB
CLIT 15.7 M 86.5 ms 9402 MiB
SRNO 0.80 M 11.2 ms 2702 MiB
MSIT 4.83 M 92.2 ms 9388 MiB
Thera Air .008 M 3.56 ms 322 MiB
Thera Plus 1.41 M 14.2 ms 664 MiB
Thera Pro 4.63 M 19.44 ms 686 MiB
Appendix G Real-World Optical Zoom Data
Table 11: Results (PSNR in dB) on the COZ test set.
@@ Appendix G Real-World Optical Zoom Data
MetaSR 28.70 26.55 25.17 24.31 23.25
LIIF 28.72 26.61 25.16 24.32 23.23
LTE 28.67 26.55 25.15 24.37 23.26
SRNO 28.73 26.59 25.15 24.31 23.25
LIT 28.74 26.58 25.15 24.35 23.19
LMI 28.86 26.66 25.22 24.39 23.29
@@
To evaluate the effectiveness of our approach on real-world continuous optical zoom data, we conducted experiments using the COZ dataset (Fu et al., 2024). We introduce Thera++, which combines two components: (1) Thera Plus (EDSR) trained on the COZ dataset, and (2) a lightweight spatial transformer network (STN) (Jaderberg et al., 2015) that estimates just 6 parameters per image to correct domain-specific affine distortions. Despite its name, the STN is not a transformer network in the modern sense that employs attention mechanisms; rather, it is an image model block that explicitly allows the spatial manipulation of data within a convolutional neural network.
Thera++ addresses an important limitation of the COZ dataset. When looking at the data it becomes obvious that COZ samples are not perfectly aligned, resulting in $xy$-jitter between images. Additionally, dynamic objects like people, dust, leaves, and moving shadows appear inconsistently across images of the same scene, significantly increasing the noise level. These challenges make super-resolution particularly difficult for this dataset. However, Thera++ outperforms previous state-of-the-art methods, including LMI, across all scaling factors, highlighting its applicability under real-world imaging conditions, see Table 11.
Appendix H Reproducibility of Existing Methods
@@
CUF (Vasconcelos et al., 2023). At the time of writing, there were no public code or checkpoints available for CUF. Therefore, we could not generate numbers for datasets and scaling factors not reported in the original paper. We have denoted those missing values with “—” in the tables. Furthermore, we could not create any qualitative samples using CUF.
CLIT (Chen et al., 2023). For CLIT, at the time of writing, there is a public code, but no checkpoints. We have made a bona fide attempt to reproduce the models, but due to the cascaded training schedule and the large model size, the training process would require excessive amounts of compute: over a month using $8\times$ Nvidia GeForce RTX 3090 GPUs. Unfortunately, the authors did not respond to our requests for the trained checkpoints used in their paper.
CiaoSR (Cao et al., 2023). We found that in the official CiaoSR implementation border cropping prior to evaluating on DIV2K deviated slightly from all other methods. We thus adapted the evaluation code to match the competition and enable a meaningful comparison. All DIV2K numbers were re-computed with the corrected code, resulting in slight deviations from those reported in the original paper.
[…]